OpenStudy (anonymous):

use an addition or subtraction formula to find the exact value of the expression tan(255degrees) So far i got tan(135+120) tan135+tan120/1-tan135tan120

OpenStudy (anonymous):

hi callisto

OpenStudy (callisto):

If you don't mind, can I introduce you a new combination of 225?

OpenStudy (callisto):

That would make life easier..

OpenStudy (anonymous):

sure

OpenStudy (anonymous):

its 255 by the way

OpenStudy (callisto):

Yes. 225 = 180 + ___?

OpenStudy (callisto):

What is that ___?

OpenStudy (anonymous):

its 255 not 225

OpenStudy (callisto):

Sorry!!!

OpenStudy (callisto):

tan(135+120) tan135+tan120/1-tan135tan120 ^Fine enough! What's the problem?

OpenStudy (anonymous):

it is tan135tan120/1-tan135tan120

OpenStudy (anonymous):

i dont know how to find tan 135 and 120

OpenStudy (callisto):

Hmmm.... \[tan(x+y) = \frac{tanx+tany}{1-tanxtany}\] So, tan135+tan120/1-tan135tan120 is fine. tan135 = tan (180 - 45) = -tan45 = ... Similar for tan120

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

now what

OpenStudy (callisto):

Can you find tan135 and tan120?

OpenStudy (anonymous):

-root 3 and -1

OpenStudy (callisto):

Plug the numbers you got into tan135+tan120/1-tan135tan120 and simplify it. Note: tan135 = -1 and tan120 = -sqrt3

OpenStudy (anonymous):